OpenStudy (anonymous):

Please help! (x-4)^2 Step by Step! WILL GIVE MEDAL!

OpenStudy (anonymous):

(x-4)^2 = (x-4)(x-4) = x(x-4) + (-4)(x-4) = x^2 - 4x - 4x + 16 = x^2 -8x + 16

OpenStudy (anonymous):

(x-4)(x-4)=(x*x-4x-4x+16) =x^2-8x+16

OpenStudy (anonymous):

OH ok!! THANKS SO MUCH!!!